how do u convert a ratio to a decimal and how do u convert each number to a fraction in lowest term
Answer by stanbon(75887)

how do u convert a ratio to a decimal?
Example 3/8=?
Divide 3 by 8 as follows:
...8)3.00000
Divide 8 into 30 to get 3 with remainder 6
Divide 8 into 60 to get 7 with remainder 40
Divide 8 into 40 to get 5 with no remainder
3/8=0.375
Comment: Not all conversions are this easy but the procedure is the same.
----------------
and how do u convert each number to a fraction in lowest term.
Example:
24/20=?
Rewrite the numerator and the denominator in factor form, as follows:
24=2*2*2*3
20=2*2*5
Then cancel the factors that are common to the numerator and denominator to get:
24/20= (2*3)/5 = 6/5
----------
Hope this helps.
Cheers,
Stan H.